OpenStudy (anonymous):

Given the two expressions shown below: square root of 64 plus square root of 5 square root of 64 plus square root of 4 Which statement best describes the two expressions? Both are rational. Both are irrational. A is rational, but B is irrational. A is irrational, but B is rational.

OpenStudy (anonymous):

@nikato

OpenStudy (anonymous):

@camerondoherty

OpenStudy (camerondoherty):

So lets find the Square root of 64 first

OpenStudy (camerondoherty):

what times what = 64?

OpenStudy (anonymous):

8

OpenStudy (camerondoherty):

yes so to simplify the question you get A\[8 + \sqrt{5}\]B\[8 + \sqrt{4}\]

OpenStudy (camerondoherty):

now what times what gives you 4?

OpenStudy (anonymous):

2

OpenStudy (camerondoherty):

Ok so B is: 8+2 what does that equal?

OpenStudy (anonymous):

10

OpenStudy (camerondoherty):

Ok so now we know that B is rational.

OpenStudy (camerondoherty):

Can you multiply 2 of the same numbers to get 5?

OpenStudy (anonymous):

Yes?

OpenStudy (camerondoherty):

Well... not exactly...

OpenStudy (camerondoherty):

\[\sqrt{5}\] stays like that which is irrational

OpenStudy (camerondoherty):

so A is irrational and B is rational

OpenStudy (camerondoherty):

Any Questions?

OpenStudy (anonymous):

This makes perfect sense! Should I tag you or just put my next question here?

OpenStudy (camerondoherty):

Well... Ask a new Question lol

OpenStudy (anonymous):

What is the value of x in the solution to the following system of equations? x − 2y = 2 3x + y = 6 Numerical Answers Expected!

OpenStudy (camerondoherty):

Are you there?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I posted the question above... ^^

OpenStudy (anonymous):

THats fine

OpenStudy (camerondoherty):

the answer is 2

OpenStudy (anonymous):

The following function represents the value of a house after x years: f(x) = 242,000(0.96)x What does 0.96 represent? The present value of the house The value of the house after x years The increase in the value of the house per year, which is 4% The decrease in the value of the house per year, which is 4%

OpenStudy (anonymous):

Cam, I closed the question. dont explain, just answer :)

OpenStudy (anonymous):

@camerondoherty ????

OpenStudy (camerondoherty):

D